The Precision Scientific Instrument Company manufactures thermometers that are supposed to give readings of 0 degrees°C

at the freezing point of water. Tests on a large sample of these thermometers reveal that at the freezing point of? water, some give readings below

0 degrees°C (denoted by negative? numbers) and some give readings above 0degrees°C

?(denoted by positive? numbers). Assume that the mean reading is

0degrees°C

and the standard deviation of the readings is

1.00degrees°C.

Also assume that the frequency distribution of errors closely resembles the normal distribution. A thermometer is randomly selected and tested. A quality control analyst wants to examine thermometers that give readings in the bottom? 4%. Find the temperature reading that separates the bottom? 4% from the others. Round to two decimal places.

A.

minus?1.48degrees°

B.

minus?1.89degrees°

C.

minus?1.63degrees°

D.

minus?1.75degrees
Answer #1

mu= 0

sigma= 1

X= unknown

probability= 0.04

z-score for probability=-1.7506

Z=(X-mu)/sigma

x=(mu+Z*sigma)

x=0-1*1.7506

x=-1.7506 degrees

D option is correct -1.75 degrees.